For the network determine the voltage Vbd in volts

Given: E1=32V; E2 =50V; R1=3200Ω; R2=5600 Ω; R3=1200 Ω;

R E Ε, RU
Answer #1

Given: -E,= 32 v E = 50v. R = 3200 R2 = 5600-2 R₂ = 1200 Ez 50V Crcuit RE =5600 si ď R =3200 S MW M & R3 =12 = 12000 Di KVL &KUL.-32 + i Riti R, -50 + {R3 = 0 i(R,+ R2 + R3)= 32 +50 ☺ 82 = 8.2.mA c 3200 +5600 + 1200 +) Cuarent is obtained Va Now volt